OpenStudy (anonymous):

What is the total number of Televisions of category-A, Refrigerators of category-B and Mobile Phones of category-A together?

OpenStudy (anonymous):

A company produces four different products, viz Mobile Phones, Televisions, Refrigerators and Computers. Each of the products has two categories, viz Category-A and Category-B. The total number of all the products is 600; 20 per cent of the total number of products is Computers and one-third of the total number of products is Televisions. 40 percent of the remaining products are Refrigerators. 50 percent of the number of Computers as well as the number Mobile Phones is of category-A. 25 per cent of the number of televisions is of category-B. One-fourth of the number of Refrigerators is of category-A.

OpenStudy (kropot72):

The number of computers of types A and B combined is 20% of 600. Can you calculate 20% of 600?

OpenStudy (anonymous):

ya i have calculated total com=120 total tele=200 total refri=240 total mobile=600 tele catb=150,comp+mob cata=120 refri cata=60 tele cata=50 mob cata=20 refri cat =180

OpenStudy (anonymous):

but i didnt get the answer

OpenStudy (kropot72):

How did you calculate the total number of refrigerators? The total of 240 is not correct.

OpenStudy (anonymous):

oh i went wrong .40/100*600.112 refri

OpenStudy (anonymous):

so 84 mobile phones

OpenStudy (kropot72):

Total refrigerators is 112. The total number of mobile phones cannot be 600. The total number of all products is 600.

OpenStudy (kropot72):

The total number of mobile phones is: 600 - (120 + 200 + 112) = ?

OpenStudy (anonymous):

84 mobile phones of cata

OpenStudy (kropot72):

Yes, there are 84 mobile phones are Category A.

OpenStudy (anonymous):

then where i gone wrong?

OpenStudy (anonymous):

314 total

OpenStudy (kropot72):

Televisions Category A are not 50.

OpenStudy (anonymous):

1/3*600=200 tele total=25/100*200 =50 cat b tele so cat a tele =200-50=150

OpenStudy (kropot72):

Correct. There are 150 televisions Category A. Please confirm your calculation for refrigerators Category B.

OpenStudy (anonymous):

k.thank u very much

OpenStudy (kropot72):

Remember that the total number of refrigerators is 112. One fourth of these are category A. So three quarters of 112 are category B refrigerators.